OpenStudy (calculusxy):

MEDAL!!!! If i'm driving in a car at 30kph and a friend is driving in a car with the same mass at 60kph, my friend's car's kinetic energy is bigger than mine by a factor of what? (that is, divide my friend's over mine.)

OpenStudy (calculusxy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Hint: Kinetic Energy (KE) KE = (1/2)*m*v^2 m = mass v = velocity

OpenStudy (calculusxy):

well since it said that the masses are the same, it would eventually cancel out. however, the velocity would be different because one is going 60 kph and the other is going 30 kph. thus, would i do: \[\frac{ 60^2 }{ 30^2 } = \frac{ 3600 }{ 900 } = 4 ? \]

jimthompson5910 (jim_thompson5910):

correct, the friend has 4 times more KE

OpenStudy (calculusxy):

i just wanted to make sure what the units would be.

OpenStudy (calculusxy):

would it just be J , like 4J

jimthompson5910 (jim_thompson5910):

The joule units would cancel though

OpenStudy (calculusxy):

only 4 right?

jimthompson5910 (jim_thompson5910):

example: your car m = 100 v = 30 friends car m = 100 v = 60 ------------------------------------------------------- Your KE KE = (1/2)*m*v^2 KE = (1/2)*100*30^2 K = 45,000 Friends KE KE = (1/2)*m*v^2 KE = (1/2)*100*60^2 K = 180,000 Ratio of the two KE values 180,000/45,000 = 4 whatever the units for the KE were, they cancel out when you divide

jimthompson5910 (jim_thompson5910):

yeah only 4

OpenStudy (calculusxy):

okay thanks :)

jimthompson5910 (jim_thompson5910):

np

OpenStudy (calculusxy):

so there is also another question on the difference on potential and kinetic energy

OpenStudy (calculusxy):

@jim_thompson5910

OpenStudy (calculusxy):

i wrote: "potential energy is the amount of energy in an object when it is still and at a certain position, however, kinetic energy is when you're calculating the amount of energy of an object when it is moving."

OpenStudy (calculusxy):

will that be fine?

jimthompson5910 (jim_thompson5910):

potential energy (PE) is the amount of stored energy that can be used and converted into KE example: pick up a rock and place it on a ledge. It has PE stored up. When you push it off the ledge, it speeds up and hits the ground converting that PE into KE

jimthompson5910 (jim_thompson5910):

potential energy due to gravity PE = m*g*h